6. In a recent poll of 755 randomly selected adults, 585 said that it is morally wrong to not report all income on tax returns. Use a 0.05 significance level to test the claim that 70% of adults say that it is morally wrong to not report all income on tax returns. Identify the null hypothesis, alternative hypothesis, test statistic, P-value, conclusion about the null hypothesis and final conclusion that addresse the original claim. Use the P-value method. Use the normal distribution as an approximation of the binomial distribution.
Identify the correct null and alternative hypotheses. Choose the correct answer below.
A. Hₒ: p = 0.3 B. Hₒ: p = 0.7
H₁: p < 0.3 H₁: P > 0.7
C. Hₒ: p = 0.7 D. Hₒ: p = 0.3
H₁: p < 0.7 H₁: p > 0.3
E. Hₒ: p = 0.7 F. Hₒ: p = 0.3
H₁: p ≠ 0.7 H₁: p ≠ 0.3

Please Some one help me with this

Answer by stanbon(75887)

6. In a recent poll of 755 randomly selected adults, 585 said that it is morally wrong to not report all income on tax returns. Use a 0.05 significance level to test the claim that 70% of adults say that it is morally wrong to not report all income on tax returns. Identify the null hypothesis, alternative hypothesis, test statistic, P-value, conclusion about the null hypothesis and final conclusion that addresse the original claim. Use the P-value method. Use the normal distribution as an approximation of the binomial distribution.
Identify the correct null and alternative hypotheses. Choose the correct answer below.
------
Note: It looks like your answer options are all confused.
The answer should be:
Ho: p = 0.7
H1: p # 0.7
---------------------
p-hat = 585/755 = 0.77
test stat: z(0.77) = (0.77-0.7)/sqrt[0.7*0.3/755] = 4.197
p-value = 2*P(z > 4.197) = 0.00002704
----
Since the p-value is less than 5%, Reject Ho.
==============================
Cheers,
Stan H.
======================
A. Hₒ: p = 0.3
B. Hₒ: p = 0.7
H₁: p < 0.3 H₁: P > 0.7
C. Hₒ: p = 0.7 D. Hₒ: p = 0.3
H₁: p < 0.7 H₁: p > 0.3
E. Hₒ: p = 0.7 F. Hₒ: p = 0.3
H₁: p ≠ 0.7 H₁: p ≠ 0.3